OpenStudy (anonymous):

what is the probability of picking a five and a jack from an ordinary deck of cards

OpenStudy (anonymous):

4 fives and 4 jacks in a deck of 52 cards

OpenStudy (anonymous):

i know so would it be (4/52)+(4/51)

OpenStudy (kropot72):

When drwing the first card the probability of either a five or a jack is 8/52. When drawing the second card the probability of drawing the required other card is 4/51. The probability of drwing afive and a jack is:\[\frac{8}{52}\times \frac{4}{51}\]

OpenStudy (anonymous):

thank you so what would be P(heart and a jack)? (16/52)(4/51)?

OpenStudy (kropot72):

There are 13 cards in the suit of hearts

OpenStudy (anonymous):

so it would be (13/52)(4/51)?

OpenStudy (kropot72):

Looks good :)

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)